# SubTrack your Grad: Gradient Subspace Tracking for Memory and Time Efficient Full-Parameter LLM Training

### v1

@@ 4 Theoretical Analysis
In this section, we analyze the convergence of SubTrack-Grad using theoretical analysis. To begin, the weights update rule of SubTrack-Grad is as follows:
$\small W_{t}=W_{0}+\sum_{t^{\prime}=0}^{t^{\prime}=t-1}\widehat{G}_{t^{\prime}}$ (7)
@@
[Convergence of SubTrack-Grad]theoremconvergence Suppose gradient has the following form (also (1)) with functions $A_{i}$, $B_{i}$, and $C_{i}$ being L-continuous as per Def. 4.1 with constants $L_{A}$, $L_{B}$, and $L_{C}$ w.r.t. weight matrix $W_{t}$; and $\|W_{t}\|_{F}\leq M$; where $W_{t}$ denotes the weight matrix at step $t$, and $M$ is a scalar value,
@@
Now, define $\widehat{B}_{i,t}=(S_{i,t}^{l})^{\top}B_{i}(W_{t})S_{i,t}^{l}$ and $\widehat{C}_{i,t}=(S_{i,t}^{r})^{\top}C_{i}(W_{t})S_{i,t}^{r}$, where $S_{i,t}^{l}$ and $S_{i,t}^{r}$ are the rank-$r$ left and right projection matrices; $B_{i}(W_{t})$ and $C_{i}(W_{t})$ denote the dependence of $B_{i}$ and $C_{i}$ on the weight matrices $W_{t}$. Further letting $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$, and $\kappa_{t}=\frac{1}{N}\sum_{i}\lambda_{min}(\widehat{B}_{i,t})\lambda_{min}(\widehat{C}_{i,t})$, where $\lambda_{min}(\cdot)$ denotes the minimum eigenvalue over each batch, and $N$ representing the number of samples in a batch. Assuming that the projection matrices remain constant during the training. Then for learning-rate $\mu$ and $min(\kappa_{t})>(L_{A}+2L_{B}L_{C}M^{2})$, the SubTrack-Grad, with $\rho_{t}\equiv 1$ (the element-wise regularizer of the optimizer) satisfies:
@@
That is, $P_{t}\rightarrow 0$ and SubTrack-Grad converges.
The proof of Theorem 4 is provided in Appendix A, based on Zhao et al. (2024). Note that while both GaLore and SubTrack-Grad assume that the subspace remains unchanged for the proof of convergence, GaLore must limit the number of updates to ensure convergence, as each update can potentially change the entire subspace. In contrast, SubTrack-Grad leverages rank-$1$ updates to the subspace, preventing drastic changes with each update. While a deeper analysis of slowly changing subspaces and their impact on convergence remains an open problem, in practice, this allows SubTrack-Grad to perform more frequent updates.
@@
The Grassmannian manifold $Gr(n,p)$ parametrizes the set of all $p$-dimensional subspaces of $\mathbb{R}^{n}$. Each point can be represented by a projection matrix $P=UU^{T}$, where $U\in St(n,p)$.
@@
[Grassmann Exponential]theoremupdaterule Let $P=UU^{T}\in\text{Gr}(n,p)$ be a point on the Grassmannian, where $U\in\text{St}(n,p)$ is the orthonormal basis of the corresponding subspace. Consider a tangent vector $\Delta\in T_{P}\text{Gr}(n,p)$, and let $\Delta_{U}^{\text{hor}}$ denote the horizontal lift of $\Delta$ to the horizontal space at $U$ in the Stiefel manifold $\text{St}(n,p)$. Suppose the thin SVD of $\Delta_{U}^{\text{hor}}$ is given by $\Delta_{U}^{\text{hor}}=\hat{Q}\Sigma V^{T},$ where $\hat{Q}\in\text{St}(n,r)$, $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the nonzero singular values of $\Delta_{U}^{\text{hor}}$ with $r=\min(p,n-p)$, and $V\in\text{St}(p,r)$. The Grassmann exponential map, representing the geodesic emanating from $P$ in the direction $\Delta$, is given by:
$\small\text{Exp}_{P}^{\text{Gr}}(t\Delta)=\big{[}UV\cos(t\Sigma)V^{T}+\hat{Q}\sin(t\Sigma)V^{T}+UV_{\perp}V_{\perp}^{T}\big{]},$
@@
This update rule generally converges to a stable subspace if the step size $\eta$ decreases over time (Balzano et al., 2011). However, a decreasing step size can impair the ability to accurately track and adapt to subspace changes. Consequently, SubTrack-Grad uses a constant step size during training and fine-tuning to effectively adjust subspaces. This approach does not hinder convergence, as proved in Theorem 4, which guarantees convergence as long as changes are controlled to maintain the stable subspace assumption.
@@ 5 Experiments and Results
Table 1: Evaluating the performance of SubTrack-Grad and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$. All hyperparameters are the same for each rank. The performance is measured via Accuracy for QNLI, MNLI, SST-2, and RTE tasks, F1 for QQP and MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA, after fine-tuning for 30 epochs.
COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg
Full-Rank 62.57 91.03 91.32 77.98 94.27 87.83 92.71 89.21 85.86
BAdam (Luo et al., 2024) 54.44 89.01 91.35 68.59 94.15 87.31 92.49 87.50 83.10
GaLore, r=4 (Zhao et al., 2024) 60.34 90.58 92.58 76.53 94.27 87.12 92.20 87.86 85.18
SubTrack-Grad, r=4 (Ours) 61.07 90.63 92.83 76.89 93.81 87.20 91.73 86.73 85.11
GaLore, r=8 (Zhao et al., 2024) 58.54 90.61 91.30 74.37 94.50 87.34 92.71 87.99 84.67
SubTrack(Last)-Grad, r=8 (Ours) 58.03 90.76 91.81 77.62 94.38 87.15 92.31 87.26 84.91
Table 2: Evaluating the performance of SubTrack-Grad and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with rank 8. All hyperparameters are consistent across different methods. The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, F1 for CB, and Matthews Correlation for AXb, after fine-tuning for 30 epochs.
BoolQ CB COPA WIC WSC AXg Avg
Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30 80.64
GaLore (Zhao et al., 2024) 85.44 88.85 80.00 71.47 63.46 100.00 81.54
BAdam (Luo et al., 2024) 82.51 53.28 59.00 70.38 60.58 51.85 62.93
SubTrack-Grad (Ours) 85.66 88.85 81.00 72.73 63.46 100.00 81.95
Figure 2: Comparing wall-time on different architecture and datasets. (a) shows the memory and wall-time measured on pretraining different Llama-architectures sizing from 60M to 3B. (b) demonstrates wall-time measured on fine-tuning RoBERTa-Large on SuperGLUE tasks. (c) presents wall-time measured on fine-tuning RoBERTa-Base on GLUE tasks. As demonstrated, GaLore can lead to 157% wall-time increase in training Llama with 3B parameters, while SubTrack-Grad significantly reduces this overhead.
To ensure a fair comparison between SubTrack-Grad and baselines, we conducted fine-tuning and pre-training experiments across various models and datasets; to measure performance, wall-time, and memory consumption, while all shared hyperparameters remain consistent.
Fine-Tuning Experiments. RoBERTa-Base and RoBERTa-Large were fine-tuned on GLUE and SuperGLUE tasks respectively; with the results presented in Table 1 and Table 2. Figures 2(b) and 2(c) compare wall-time during fine-tuning, with detailed reports provided in Appendix D, as demonstrated, SubTrack-Grad enjoys the same memory reduction as GaLore. For comparing wall-times, models were fine-tuned up to a number of iterations to exactly include five subspace updates, while the evaluation steps were excluded to provide an accurate comparison. SubTrack-Grad demonstrates significant efficiency improvements compared to GaLore by reducing wall-time by up to 20.57% on GLUE tasks (15% average reduction) and up to 65% on SuperGLUE tasks (22% average reduction). Despite limiting updates to rank-$1$ modifications of the previous subspace, SubTrack-Grad achieves superior or comparable performance compared to GaLore. In addition, BAdam is a computationally efficient method due to optimizing one block at a time; but its inability to update all the parameters simultaneously has caused performance loss, and the loss gets greater as the model size increases. Details of fine-tuning hyperparameters are provided in Appendix E.
Pre-Training Experiments. Different Llama-based architectures were pre-trained on C4 datasets for 10K steps, and the results are reported in Table 3. For wall-time comparison, each architecture was trained for 2K iterations, with the subspace update interval set to 200, ensuring exactly 10 subspace updates. Figure 2(a) illustrates wall-times, with detailed reports provided in Appendix D that indicates same memory consumption between GaLore and SubTrack-Grad. Notably, GaLore incurred a substantial 157% increase in wall-time compared to full-rank training for the model of size 3B, whereas SubTrack-Grad exhibited only a 31% increase—representing a 49% reduction in wall-time compared to GaLore. On average, SubTrack-Grad required 27% less training time across these models. Additionally, BAdam demonstrate significant runtime reduction, but its poor performance in pre-training experiments due to its partial optimization, makes it an unsuitable choice for training from scratch. Additional architectural details for the Llama-based models are provided in Appendix F.
Table 3: Evaluation loss comparison for pre-training different Llama-based architectures on C4 dataset after 10K iterations.
60M 130M 350M 1B 3B Avg
r=128 r=256 r=256 r=512 r=512 -
Full-Rank 3.41 3.25 3.40 4.61 4.52 3.84
GaLore 4.02 3.61 3.62 6.53 6.57 4.87
BAdam 7.86 7.08 7.62 7.28 7.12 7.39
SubTrack-Grad 4.13 3.77 3.82 6.53 6.59 4.97
@@
Time and Space Complexity. Table 4 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$. As discussed earlier, GaLore performs SVD on the gradient matrix to estimate the underlying subspace, whereas SubTrack-Grad uses subspace tracking. Comparing the time complexities of these methods, highlights why SubTrack-Grad is significantly more efficient than GaLore. A breakdown of the subspace update time complexity for SubTrack is shown in Appendix C. As shown in Figure 2(a), for pre-training a Llama model of size 3B, GaLore incurs over 157% wall-time overhead, compared to 31% for SubTrack-Grad. This demonstrates a 49% wall-time reduction compared to GaLore. Additionally, the memory required for storing optimizer states in SubTrack-Grad, is equivalent to GaLore’s.
Table 4: Optimizer’s memory requirement in each method along with the time complexity of involved subspace update operations, considering a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$.
Optimizer Memory Subspace Update Time Complexity
@@
GaLore $mr+2nr$ $O(nm^{2})$
SubTrack-Grad $mr+2nr$ $O(mnr)$
Reducing Optimization Jumps. We argue that GaLore is susceptible to noise, as it relies on the SVD of a single gradient matrix at each subspace update step. Figure 3 compares SubTrack-Grad and GaLore on the Ackley function to illustrate potential jumps and deviations. These jumps prevent GaLore from reaching the global minimum within 100 steps when the scale factor is set to 1. While the scale factor to 3 enables GaLore to reach that minimum, it also results in larger jumps. Figure 5 in Appendix H illustrates that the overall convergence during both pre-training and fine-tuning is similar to other methods when using SubTrack-Grad.
Figure 3: Comparison of the Adam Optimizer Combined with SubTrack-Grad (a, d) and GaLore (b, e) on the Ackley Function for 100 optimization steps and the subspace update interval set to 10. With scale factor 1, GaLore can not reach the to the global minimum of the function due to its abrupt jumps, and for a scale factor of 3, the length of jumps increase. The 3D plots (c, f) provide visualization of the dynamics of the Ackley function and the optimization paths followed by each method in a single plot.
Runtime Consistency. Figure 4 compares the performance and wall-times of GaLore and SubTrack-Grad across subspace update intervals ranging from 50 to 500 while fine-tuning RoBERTa-Base on the COLA task and pre-training a Llama architecture with 60M parameters on the C4 dataset. The subspace update interval denotes the number of iterations between two updates; thus, increasing this interval reduces the update frequency. GaLore’s runtime significantly increases with more frequent subspace updates, whereas SubTrack-Grad maintains minimal overhead. In Table 15 different performance achieved on COLA dataset for different subspace update intervals are compared, illustrating that increasing the update frequency significantly raises GaLore’s runtime, while the performance of the two methods remains very similar.
Figure 4: Comparing runtimes of GaLore and SubTrack-Grad. RoBERTa-Base is fine-tuned for 10 epochs on COLA and Llama-60M is pre-trained for 2500 iterations on C4 dataset. Subspace update intervals range from 50 to 500 in both experiments. Notice that by increasing subspace update interval, the update frequency actually decreases, as it indicates the number of iterations between two subspace update steps. As demonstrated, GaLore’s wall-time can increase drastically when update frequency increases.

### v2

@@ 3 Theoretical Analysis
In this section, we analyze the convergence of Grassmannian Subspace Tracking applied in SubTrack++ using theoretical analysis. To begin, the general weights update rule is as follows:
$\small W_{t}=W_{0}-\alpha\cdot\sum_{t^{\prime}=0}^{t^{\prime}=t-1}\widehat{G}_{t^{\prime}}$ (13)
@@
[Convergence of Grassmannian Subspace Tracking]theoremconvergence Suppose gradient has the following form with functions $A_{i}$, $B_{i}$, and $C_{i}$ being L-continuous as per Def. 3.1 with constants $L_{A}$, $L_{B}$, and $L_{C}$ w.r.t. weight matrix $W_{t}$; and $\|W_{t}\|_{F}\leq M$; where $W_{t}$ denotes the weight matrix at step $t$, and $M$ is a scalar value,
@@
Now, define $\widehat{B}_{i,t}=(S_{i,t}^{l})^{\top}B_{i}(W_{t})S_{i,t}^{l}$ and $\widehat{C}_{i,t}=(S_{i,t}^{r})^{\top}C_{i}(W_{t})S_{i,t}^{r}$, where $S_{i,t}^{l}$ and $S_{i,t}^{r}$ are the rank-$r$ left and right projection matrices; $B_{i}(W_{t})$ and $C_{i}(W_{t})$ denote the dependence of $B_{i}$ and $C_{i}$ on the weight matrices $W_{t}$. Further letting $P_{t}={S_{t}^{l}}^{\top}G_{t}S_{t}^{r}$, and $\kappa_{t}=\frac{1}{N}\sum_{i}\lambda_{min}(\widehat{B}_{i,t})\lambda_{min}(\widehat{C}_{i,t})$, where $\lambda_{min}(\cdot)$ denotes the minimum eigenvalue over each batch, and $N$ representing the number of samples in a batch. Assuming that the projection matrices remain constant during the training. Then for learning-rate $\mu$ and $min(\kappa_{t})>(L_{A}+2L_{B}L_{C}M^{2})$, subspace tracking, with $\rho_{t}\equiv 1$ (the element-wise regularizer of the optimizer) satisfies:
@@
That is, $P_{t}\rightarrow 0$ and it converges.
The proof of Theorem 3 is provided in Appendix A, based on Zhao et al. (2024). While both GaLore and SubTrack++ assume the subspace remains unchanged for the proof of convergence, GaLore must limit these updates to ensure convergence, as each update can potentially change the entire subspace. In contrast, SubTrack++ leverages rank-$1$ updates to the subspace, preventing drastic changes with each update. While a deeper analysis of slowly changing subspaces and their impact on convergence remains an open problem, in practice, this allows SubTrack++ to perform more frequent updates.
@@
The Grassmannian manifold $Gr(n,p)$ parametrizes the set of all $p$-dimensional subspaces of $\mathbb{R}^{n}$. Each point can be represented by a projection matrix $P=UU^{\top}$, where $U\in St(n,p)$.
@@
[Grassmann Exponential]theoremupdaterule Let $P=UU^{\top}\in\text{Gr}(n,p)$ be a point on the Grassmannian, where $U\in\text{St}(n,p)$ is the orthonormal basis of the corresponding subspace. Consider a tangent vector $\Delta\in T_{P}\text{Gr}(n,p)$, and let $\Delta_{U}^{\text{hor}}$ denote the horizontal lift of $\Delta$ to the horizontal space at $U$ in the Stiefel manifold $\text{St}(n,p)$. Suppose the thin SVD of $\Delta_{U}^{\text{hor}}$ is given by $\Delta_{U}^{\text{hor}}=\hat{Q}\Sigma V^{\top},$ where $\hat{Q}\in\text{St}(n,r)$, $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the nonzero singular values of $\Delta_{U}^{\text{hor}}$ with $r=\min(p,n-p)$, and $V\in\text{St}(p,r)$. The Grassmann exponential map, representing the geodesic emanating from $P$ in the direction $\Delta$, is given by:
$\small\text{Exp}_{P}^{\text{Gr}}(t\Delta)=\big{[}UV\cos(t\Sigma)V^{\top}+\hat{Q}\sin(t\Sigma)V^{\top}+UV_{\perp}V_{\perp}^{\top}\big{]},$
@@
This update rule generally converges to a stable subspace if the step size $\eta$ decreases over time (Balzano et al., 2011). However, a decreasing step size can impair the ability to accurately track and adapt to subspace changes. Consequently, SubTrack++ uses a constant step size to effectively adjust subspaces. This approach does not hinder convergence, as proved in Theorem 3, which guarantees convergence as long as changes are controlled to maintain the stable subspace assumption.
@@ 4 Experiments and Results
To ensure fair comparison, we evaluated SubTrack++ across diverse models and datasets through pre-training and fine-tuning, measuring key metrics critical to LLM democratization.
Pre-Training Experiments. We pre-trained several Llama-based models on the C4 dataset, with results in Table 1. To ensure a fair comparison, we benchmarked against a diverse set of baselines. BAdam (Luo et al., 2024) is notably efficient in both memory and time, but its reliance on partial parameter updates leads to reduced performance. GaLore (Zhao et al., 2024) supports full-parameter training with low memory usage, though this comes at the cost of more than 150% increase in wall-time relative to AdamW in large models.
Table 1: We compare evaluation loss for pre-training Llama-based architectures on the C4 dataset over 10K iterations. SubTrack++ outperforms all other baselines in nearly every configuration. The best results are marked in bold, with the second-best performance underlined. LDAdam could not be run on the 7B configuration due to an out-of-memory error with our available resources.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 3.41 3.25 3.40 4.61 4.52 4.30
GaLore (Zhao et al., 2024) 4.02 3.61 3.62 6.53 6.57 5.55
BAdam (Luo et al., 2024) 7.86 7.08 7.62 7.28 7.12 6.76
Online Subspace Descent (Liang et al., 2024) 4.18 3.88 4.09 6.79 6.85 5.69
LDAdam (Robert et al., 2025) 3.52 3.44 3.67 4.70 4.39 OOM
Fira (Chen et al., 2025) 3.80 3.55 3.56 6.31 6.50 6.83
SubTrack++ (Ours) 3.43 3.24 3.29 4.51 4.49 4.62
As another relative baseline, Online Subspace Descent (Liang et al., 2024) uses Online PCA to track gradient subspace. LDAdam (Robert et al., 2025) and Fira (Chen et al., 2025) deliver strong performance at the cost of increasing wall-time even more than GaLore. Across these methods, SubTrack++ achieves SOTA results while using 14% less memory and running 43% faster than LDAdam, the most accurate baseline on 1B model. Hyperparameter details are provided in Appendix G, with a detailed wall-time and memory analysis in Appendix F.
Figure 3: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ . While Grassmannian tracking alone matches GaLore’s step-wise convergence, it significantly reduces wall-time, as shown in Figure LABEL:fig:modular.
Figure LABEL:fig:modular compares pre-training performance of a 1B-parameter Llama model across several baselines. SubTrack++ converges fastest in both steps and wall-time, emphasizing the power of geometry-aware optimization in boosting performance, reducing resource demands, and democratizing access to SOTA training methods.
Ablation Study. We conducted an ablation study to evaluate the individual and combined effects of projection-aware optimizers and recovery scaling, integrated with Grassmannian subspace tracking on the 1B Llama model. We use the experimental settings outlined in Table 10. As shown in Figure 3, projection-aware optimizer and recovery scaling independently yield substantial improvements over pure Grassmannian subspace tracking, reducing loss from 6.53 to 5.43 and 5.28, respectively; while their combination, SubTrack++ , achieves 4.51 surpassing all baselines. Notably, these gains are achieved with minimal increases in runtime and memory.
Fine-Tuning Experiments. RoBERTa-Base and RoBERTa-Large are fine-tuned on GLUE and SuperGLUE tasks; with the results presented in Table 4 and 5, respectively. More details and hyperparameters are provided in Appendix E.
@@
Time and Space Complexity. Table 2 provides memory requirements of the optimizer states and the time complexity of the subspace update step considering an $m\times n$ gradient matrix with $m\leq n$. GaLore (Zhao et al., 2024) and Fira (Chen et al., 2025) periodically perform SVD to estimate the underlying subspace, while LDAdam (Robert et al., 2025) relies on the faster PowerSGD to update the subspace at every iteration. In contrast, SubTrack++ employs Grassmannian-based subspace tracking at the same frequency as GaLore and Fira. Comparing the time complexities of these methods, highlights why SubTrack++ is significantly more efficient than SVD-based methods. A breakdown of the subspace update time complexity for SubTrack++ is shown in Appendix D. Additionally, the memory required for storing optimizer states in SubTrack++ , is equivalent to GaLore and other baselines.
Table 2: The optimizer’s state parameter count and subspace update time complexity across baselines, given a gradient matrix of dimension $m\times n$ and projection rank $r$ where $r\ll m\leq n$. LDAdam updates the subspace at every iteration, while other methods update it every $k$ steps.
Optimizer Mem. Subspace Update Time
@@
LDAdam $mr+2nr$ $O(mnr)$
GaLore, Fira $mr+2nr$ $O(nm^{2})$
SubTrack++ $mr+2nr$ $O(mnr)$
(a) Ours, SF=1
(b) GaLore, SF=1
(c) Ours, SF=3
(d) GaLore, SF=3
Figure 5: Comparison of Grassmannian subspace tracking (Ours) (a, c) and GaLore’s SVD (b, d) on the Ackley Function over 100 optimization steps, with a subspace update interval of 10. SF stands for scale factor; with a scale factor of 1, GaLore fails to reach the global minimum due to abrupt jumps. At a scale factor of 3, while the minimum is reached, the jump length increases. This demonstrates SVD’s sensitivity to noise and abrupt changes, highlighting the robustness of our Subspace Tracking method with its controlled subspace updates.
Robust Subspace Tracking. Relying on SVD for subspace updates makes methods sensitive to noise and abrupt changes (He et al., 2025). Figure 5 compares Grassmannian subspace tracking with GaLore’s SVD on the Ackley function, highlighting how SVD causes erratic jumps, while our subspace tracking ensures robust optimization. GaLore struggles to reach the global minimum within 100 steps at scale factor 1, and although increasing the scale factor to 3 improves performance, it amplifies jumps that hinder convergence in non-convex settings, revealing sensitivity to hyperparameters, noise, and abrupt changes.
5 Related Works
Parameter-Efficient Training. Several works aim to improve the efficiency of training LLMs, addressing a growing demand as their popularity rapidly increases. Popular LoRA (Hu et al., 2021) significantly reduces memory requirements for fine-tuning LLMs by leveraging two low-rank trainable low-rank matrices. Dettmers et al. (2024) employ quantization techniques and paged optimizers to further reduce memory usage. Additionally, Yaras et al. (2024) introduce Deep LoRA to address overfitting issues and reducing the need for precise tuning of the rank parameter. Several other works have also extended LoRA to enhance the efficiency of training and fine-tuning LLMs (Lialin et al., 2023; Renduchintala et al., 2024; Xia et al., 2024; Pan et al., 2024). Miles et al. (2024) propose compressing intermediate activations and reconstructing them during backpropagation to enhance memory efficiency. Additionally, Hao et al. (2024) demonstrate that full-parameter fine-tuning is feasible by random projections on the gradient matrix, showing that LoRA essentially performs a down-projection of the gradient. BAdam (Luo et al., 2024) leverages the block coordinate descent framework to reduce memory consumption while maintaining capabilities comparable to Adam.
Gradient Low-Rank Projection. Several approaches aim to reduce optimizer states, as optimizers like Adam (Kingma and Ba, 2017) account for a significant portion of memory footprint (Li et al., 2023; Anil et al., 2019; Lv et al., 2024; Dettmers et al., 2022). MicroAdam (Modoranu et al., 2024) tackles this by compressing the gradient space and utilizing the compression error through feedback loops. Adam-mini (Zhang et al., 2024) partitions model into blocks, assigning a single learning rate to each block to preserve performance while saving memory. Gur-Ari et al. (2018); Schneider et al. (2024); Yaras et al. (2023) suggest that a substantial portion of gradients lies within a largely consistent subspace. GaLore (Zhao et al., 2024) leverages this fact to reduce the optimizer’s memory by projecting gradients into a low-rank subspace and then projecting them back for full parameter tuning. This approach has been integrated with other methods regarding efficient LLM training (Li et al., 2024; Rajabi and Rambhatla, 2024b). However, not all layers’ gradients evolve within a stable low-rank subspace. Jaiswal et al. (2024); Rajabi and Rambhatla (2024d) identify layers where gradients evolve within a low-dimensional subspace and fine-tune only those layers, freezing the others to avoid inefficient low-rank updates. Grass (Muhamed et al., 2024) reduces memory usage by applying sparse projection matrices to the gradient. Ramesh et al. (2024) dynamically select and update a subset of parameters, for a fast and memory-efficient training. FIRA (Chen et al., 2025) utilize a norm-based scaling method along with GaLore to maintain performance comparable to full-rank training. GoLore (He et al., 2025) addresses GaLore’s convergence issues and employ random projection in latter steps as a solutions. LDAdam (Robert et al., 2025) performs optimization within lower-dimensional subspaces, incorporating a projection-aware optimization update rule and a generalized error feedback mechanism. APOLLO (Zhu et al., 2025) approximates channel-wise learning rate scaling based on random projection. Also, Liang et al. (2024) introduce a dynamically evolving projection matrix updated via online PCA, enhancing the model’s ability to navigate the parameter space efficiently without relying on expensive SVD, a similar approach is also proposed in Rajabi and Rambhatla (2024a).
Grassmannian and Subspace Tracking. A common approach in working with high-dimensional data is to project the data into a lower-dimensional space, and many studies focus on tracking these subspaces as they evolve. Balzano et al. (2011); Rajabi and Rambhatla (2024c) introduce an incremental method for updating subspaces on the Grassmannian when the data is partially observed. Zhang and Balzano (2016) and Kasai (2017) propose methods to handle noise effect in tracking these subspaces. Furthermore, Blocker et al. (2023) present a method for evolving geodesic-based data in the Grassmannian for updating the subspace effectively.